OpenStudy (anonymous):

Graph this parametric pair: x = 7 sin t + sin 7t y = 7 cos t + cos 7t

OpenStudy (anonymous):

I'll start by plugging in the same values as before; 0, pi/4, pi/2, etc?

OpenStudy (anonymous):

Go ahead with these values.

OpenStudy (anonymous):

x = 7 sin (pi) + sin (7pi) = -4E-13 What did I do worng haha?

OpenStudy (anonymous):

t = pi, x = 0

OpenStudy (anonymous):

sin n*pi = 0 for n is an integer

OpenStudy (anonymous):

But what do I do about the sin 7t?

OpenStudy (anonymous):

t = pi, sin 7t = sin 7pi = sin (3*2pi + pi) = sin pi = 0

OpenStudy (anonymous):

The graph of this one looks crazy, I don't think I need to do this by hand, maybe I can just use a program.

OpenStudy (anonymous):

You may use a program or I would suggest you to include t = pi/6 and pi/3 also in the first quadrant.Plot it in first quadrant. What you have got in first quadrant, just repeat it in other quadrants symmetrically

OpenStudy (anonymous):

You don't need to do calculations for other quadrants.

OpenStudy (anonymous):

Alright. I'll just do the first quadrant for this. Thanks for your help.

OpenStudy (anonymous):

You are welcome :)

OpenStudy (anonymous):

If i have more questions will you still be available haha?